# M2DAO-Talker: Harmonizing Multi-granular Motion Decoupling and Alternating Optimization for Talking-head Generation

### v2

@@ Abstract
Audio-driven talking head generation holds significant potential for film production. While existing 3D methods have advanced motion modeling and content synthesis, they often produce rendering artifacts, such as motion blur, temporal jitter, and local penetration, due to limitations in representing stable, fine-grained motion fields. Through systematic analysis, we reformulate talking head generation into a unified framework comprising three steps: video preprocessing, motion representation, and rendering reconstruction. This framework underpins our proposed M2DAO-Talker, which addresses current limitations via multi-granular motion decoupling and alternating optimization. Specifically, we devise a novel 2D portrait preprocessing pipeline to extract frame-wise deformation control conditions (motion region segmentation masks, and camera parameters) to facilitate motion representation. To ameliorate motion modeling, we elaborate a multi-granular motion decoupling strategy, which independently models non-rigid (oral and facial) and rigid (head) motions for improved reconstruction accuracy. Meanwhile, a motion consistency constraint is developed to ensure head-torso kinematic consistency, thereby mitigating penetration artifacts caused by motion aliasing. In addition, an alternating optimization strategy is designed to iteratively refine facial and oral motion parameters, enabling more realistic video generation. Experiments across multiple datasets show that M2DAO-Talker achieves state-of-the-art performance, with the 2.43 dB PSNR improvement in generation quality and 0.64 gain in user-evaluated video realness versus TalkingGaussian while with 150 FPS inference speed. Our project homepage is https://m2dao-talker.github.io/M2DAO-Talk.github.io
@@ Introduction
Figure 2: Detailed architecture of the network designed to capture non-rigid facial dynamics. “SA” stands for Self-Attention module.
@@
Dataset. Following prior works (Ye et al. 2023; Li et al. 2023; Guo et al. 2021; Tang et al. 2022), we use seven publicly available video clips to train and evaluate both our M2DAO-Talker and baselines. Each video averages 7,066 frames at 25 FPS, focusing on a centered portrait, which are $512\times 512$ (“Obama2”,“May”, “Shaheen”and “Lieu”) or $450\times 450$ (“Obama”, “Jae-in”, and “Obama1”) resolution.
@@
Method “Shaheen” Audio “Lieu” Audio
“Obama” “May” “Obama” “May”
@@
Figure 4: Qualitative results of Image Quality Comparison. Our method effectively eliminates edge blur and corrects facial distortions, blinking errors, and unnatural head-torso synthesis that were present in previous approaches. For improved visualization, please zoom in.
@@
Ablation Setting. To assess the contribution of each module, we conduct ablation studies on key components across our framework. For MRS and HPE, which involve specific backbone design choices, we evaluate by replacing them with widely adopted alternatives under controlled settings. Specifically, we substitute MRS with BiseNet (Yu et al. 2018) and conduct experiments on datasets such as “Jae-in”,“Shaheen”, and “May” where boundary ambiguity is prominent. For HPE, we revert to 3DMM-based camera estimation and replace other methods’ estimations with HPE. We evaluate the results on the “Obama”, “Obama1”, and “Obama2” sequences, which feature diverse head movements. In contrast, we ablate AOS and MCC on all seven datasets to comprehensively assess the global impact.
Backbone MRS BN PSNR $\uparrow$ LPIPS $\downarrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 35.041 0.0223 2.4575
34.400 0.0243 2.4586
TalkingGaussian (Li et al. 2025) 31.519 0.0360 2.5973
33.260 0.0291 2.5603
SyncTalk (Peng et al. 2024) 33.955 0.0249 2.5374
34.355 0.0261 2.5733
@@
Ablation Results. Replacing the standard segmentation backbone with MRS yields consistent improvements (Table 4). These gains stem from MRS’s precise boundary localization, which reduces motion interference and enables more accurate region-specific deformation. Even SyncTalk, which lacks explicit deformation modeling, benefits from MRS, confirming its general utility for motion-aware synthesis. Incorporating HPE also improves all metrics (Table 5), as it explicitly decouples rigid head rotation from facial dynamics, enabling smoother and more stable deformation learning, especially under large pose variations. With HPE, the attention maps of upper-face ($V_{e}$) and lower-face ($V_{a}$) focus more on the corresponding regions (Figure 6), indicating that accurate decoupling of rigid head rotation enhances the learning of non-rigid facial motion.
@@
Backbone HPE PSNR $\uparrow$ LPIPS $\downarrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 33.746 0.0249 2.809
33.648 0.0249 2.799
TalkingGaussian (Li et al. 2025) 32.384 0.0294 2.825
33.215 0.0265 2.762
SyncTalk (Peng et al. 2024) 33.439 0.0260 2.839
33.669 0.0260 2.787
@@
Backbone MCC AOS PSNR $\uparrow$ LPIPS $\downarrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 34.475 0.0229 2.636
34.123 0.0239 2.624
34.375 0.0229 2.620
@@ Conclusion
  • D. Additional Visualization. We present qualitative comparisons against SyncTalk and TalkingGaussian on high-resolution real-world data. We further analyze common failure modes to highlight our model’s advantages in robustness and realism.
@@
However, existing 3DGS-based approaches primarily derive motion dynamics function from temporal sequences, where deformation fields are conditioned solely on frame indices or timestamps. Previous works reveal that nonverbal facial actions introduce motion variances uncorrelated with phoneme timing and features alter articulation dynamics beyond temporal alignment. It indicates that audio-driven motions are governed with multiple factors where time serves merely as a carrier variable. To bridge this gap, we extract phoneme-aligned mouth movements features and person-agnostic facial expression features from 2D video as control conditions. And we establish a nonlinear mapping from control conditions to Gaussian’s deformation, synthesizing motions congruent with both audio content and expressive context.
@@
Audio-Visual Encoder Talking head generation relies on strong correlations between phonemes and orofacial kinematics. While existing audio encoders (e.g., automatic speech recognition (ASR)-focused models) prioritize semantic extraction of speech signals, they fail to capture fine-grained mouth movement features critical for articulatory dynamics. To address this, we adopt SyncTalk’s pre-trained audio-visual encoder (Peng et al. 2024), adversarially trained on the LRS2 dataset (Afouras et al. 2018) with a lip-sync discriminator. This model extracts phoneme-aware mouth movement features $a$, establishing cross-modal alignment between speech signals and articulatory motion. These features are injected into the deformation field via attention-based modulation, controlling lip-related Gaussian primitive deformations.
Facial Action Encoder Facial expressions during speech production involve complex muscle activations that extend beyond phoneme articulation, including subtle movements such as squinting, eyebrow elevation, and frowning. Current approaches exhibit limitations in modeling these dynamics. For example, some efforts regulate basic blinking patterns (Li et al. 2023; Tang et al. 2022; Liu and Hao 2023) but fail to capture nuanced facial muscle coordination. Blendshape-driven techniques (Peng et al. 2024) enable broader expression control through 3D shape approximations but lack direct correspondence to biomechanical muscle activations. To bridge this gap, we propose a facial action encoder grounded in the Facial Action Coding System (FACS) (Ekman and Friesen 1978). Our framework parameterizes facial dynamics using six anatomically defined Action Units (AUs: 1, 4, 5, 6, 7, 45), selected for two key advantages.
@@
Audio-Visual Encoder. To assess the effectiveness of AVE, we conduct ablation experiments on datasets with constrained head rotations (“Lieu”, “Shaheen” and “Jae-in”), where subtle articulatory dynamics are critical for accurate lip synchronization. Specifically, for AVE-equipped methods (our M2DAO-Talker approach and SyncTalk (Peng et al. 2024)): we replace AVE with DeepSpeech (DS) audio encoding. For DS-based methods (TalkingGaussian (Li et al. 2025)), we substitute DS with AVE encoding. We maintain the identical architectures otherwise for fair comparison, and evaluate their performance across PSNR, LPIPS, and LMD. Experiments in Table 5 demonstrate that AVE implementations outperform DS variants across all metrics, which displays AVE’s advantage in capturing mouth movement dynamics over DS’s semantic speech features. It is speculated that AVE’s focus on articulatory features enhances visual-articulatory alignment. Significantly, our M2DAO-Talker method maintains performance parity with SyncTalk and TalkingGaussian even when using suboptimal DS encoding, demonstrating obvious architectural resilience and superiority.
Backbone AVE DS PSNR $\uparrow$ LPIPS $\downarrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 35.932 0.0164 2.423
35.672 0.0167 2.484
TalkingGaussian (Li et al. 2025) 32.687 0.0246 2.484
32.841 0.0242 2.363
SyncTalk (Peng et al. 2024) 35.075 0.0178 2.503
34.878 0.0181 2.510
@@
Backbone AU BS PSNR $\uparrow$ LPIPS $\downarrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 34.475 0.0229 2.636
34.333 0.0226 2.672
TalkingGaussian (Li et al. 2025) 32.042 0.0319 2.712
31.938 0.0320 2.712
SyncTalk (Peng et al. 2024) 33.902 0.0246 2.685
33.905 0.0252 2.660
@@
Facial Action Encoder. To evaluate the efficacy of FAE, we conduct cross-paradigm ablation experiments by substituting Action Units (AUs) and Blendshapes (BS) across seven datasets with varying expression complexity. For AU-based methods (our M2DAO-Talker approach and TalkingGaussian (Li et al. 2025)), we replace AUs with BS. Conversely, for BS-based methods, we substitute BS with AUs for facial expression control. The performance comparison across PSNR, LPIPS, and LMD are shown in Table 6, revealing that AU-to-BS substitution causes consistent metric degradation on average (PSNR: --0.142 dB, LMD: $+$0.036). Notably, SyncTalk exhibits atypical behavior: BS substitution improves LPIPS by 2.43% despite minor PSNR/LMD changes. It is supposed that BS’s holistic facial parameterization better aligns with SyncTalk’s motion field paradigm. Overall, AUs outperforms BS in almost scenarios, particularly in complex and diverse expression scenarios (e.g., “May”:$+$0.61 dB, “Obama1” :$+$0.29 dB and “Obama2” :$+$0.19 dB). This validates AUs’ superiority in modeling fine-grained muscle activations critical for subtle expressions.
@@
To facilitate a more comprehensive comparison with SyncTalk and TalkingGaussian, we select five speech videos exceeding four minutes in duration from the High-Definition Talking Face (HDTF) dataset. Unlike the datasets used in the main paper, these videos feature substantial torso motion and large-scale head rotations during speech, thereby presenting challenging real-world conditions that stress model generalization and robustness. We adopt the same self-reconstruction setup as in earlier experiments, and visualize the results in Figure 7. The following analysis highlights qualitative differences across the three methods.

### v3

@@ Abstract
Audio-driven talking head generation holds significant potential for film production. While existing 3D methods have advanced motion modeling and content synthesis, they often produce rendering artifacts, such as motion blur, temporal jitter, and local penetration, due to limitations in representing stable, fine-grained motion fields. Through systematic analysis, we reformulate talking head generation into a unified framework comprising three steps: video preprocessing, motion representation, and rendering reconstruction. This framework underpins our proposed M2DAO-Talker, which addresses current limitations via multi-granular motion decoupling and alternating optimization. Specifically, we devise a novel 2D portrait preprocessing pipeline to extract frame-wise deformation control conditions (motion region segmentation masks, and camera parameters) to facilitate motion representation. To ameliorate motion modeling, we elaborate a multi-granular motion decoupling strategy, which independently models non-rigid (oral and facial) and rigid (head) motions for improved reconstruction accuracy. Meanwhile, a motion consistency constraint is developed to ensure head-torso kinematic consistency, thereby mitigating penetration artifacts caused by motion aliasing. In addition, an alternating optimization strategy is designed to iteratively refine facial and oral motion parameters, enabling more realistic video generation. Experiments across multiple datasets show that M2DAO-Talker achieves state-of-the-art performance, with the 2.43 dB PSNR improvement in generation quality and 0.64 gain in user-evaluated video realness versus TalkingGaussian while with 150 FPS inference speed.
@@ Introduction
Figure 2: Detailed architecture of the network designed to capture non-rigid facial dynamics. ``SA'' stands for Self-Attention module.
@@
Dataset. Following prior works (Ye et al. 2023; Li et al. 2023; Guo et al. 2021; Tang et al. 2022), we use seven publicly available video clips to train and evaluate both our M2DAO-Talker and baselines. Each video averages 7,066 frames at 25 FPS, focusing on a centered portrait, which are $\times 512$ 512 (``Obama2'',``May'', ``Shaheen''and ``Lieu'') or $\times 450$ 450 (``Obama'', ``Jae-in'', and ``Obama1'') resolution.
@@
Method ``Shaheen" Audio ``Lieu" Audio
``Obama" ``May" ``Obama" ``May"
@@
Figure 4: Qualitative results of Image Quality Comparison. Our method effectively eliminates edge blur and corrects facial distortions, blinking errors, and unnatural head-torso synthesis that were present in previous approaches. For improved visualization, please zoom in or refer to Supplementary Material.
@@
Ablation Setting. To assess the contribution of each module, we conduct ablation studies on key components across our framework. For MRS and HPE, which involve specific backbone design choices, we evaluate by replacing them with widely adopted alternatives under controlled settings. Specifically, we substitute MRS with BiseNet (Yu et al. 2018) and conduct experiments on datasets such as ``Jae-in'',``Shaheen'', and ``May'' where boundary ambiguity is prominent. For HPE, we remove the high flow error frames filter and replace the other methods' estimations with HPE. We evaluate the results on the “Obama”, “Obama1”, and “Obama2” sequences, which feature diverse head movements. In contrast, we ablate AOS and MCC on all seven datasets to comprehensively assess the global impact.
Backbone MRS BN PSNR $\uparrow$ Sync-C $\uparrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 35.041 7.965 2.4575
34.400 7.956 2.4586
TalkingGaussian (Li et al. 2025) 31.519 5.971 2.5973
33.260 6.868 2.5603
SyncTalk (Peng et al. 2024) 33.955 7.663 2.5374
34.355 7.819 2.5733
@@
Ablation Results. Replacing the standard segmentation backbone with MRS yields consistent improvements (Table 4). These gains stem from MRS’s precise boundary localization, which reduces motion interference and enables more accurate region-specific deformation. Even SyncTalk, which lacks explicit deformation modeling, benefits from MRS, confirming its general utility for motion-aware synthesis. Incorporating HPE also improves most metrics (Table 5), as it explicitly decouples rigid head rotation from facial dynamics, enabling smoother and more stable deformation learning, especially under large pose variations. With HPE, the attention maps of upper-face ($V_{e}$) and lower-face ($V_{a}$) focus more on the corresponding regions (Figure 6), indicating that accurate decoupling of rigid head rotation enhances the learning of non-rigid facial motion.
@@
Backbone HPE PSNR $\uparrow$ Sync-C $\uparrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 33.746 7.372 2.809
33.648 7.338 2.799
TalkingGaussian (Li et al. 2025) 32.384 6.836 2.825
33.215 6.859 2.762
SyncTalk (Peng et al. 2024) 33.439 7.450 2.839
33.669 7.174 2.787
@@
Backbone MCC AOS PSNR $\uparrow$ Sync-C $\uparrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 34.475 7.757 2.636
34.123 7.688 2.624
34.375 7.556 2.620
@@ Conclusion
  • D. Additional Visualization. We present qualitative comparisons against SyncTalk and TalkingGaussian on high-resolution real-world data. We further analyze common failure modes to highlight our model's advantages in robustness and realism.
@@
However, existing 3DGS-based approaches primarily derive motion dynamics function from temporal sequences, where deformation fields are conditioned solely on frame indices or timestamps. Previous works reveal that nonverbal facial actions introduce motion variances uncorrelated with phoneme timing and features alter articulation dynamics beyond temporal alignment. It indicates that audio-driven motions are governed with multiple factors where time serves merely as a carrier variable. To bridge this gap, we extract phoneme-aligned mouth movements features and person-agnostic facial expression features from 2D video as control conditions. And we establish a nonlinear mapping from control conditions to Gaussian's deformation, synthesizing motions congruent with both audio content and expressive context.
@@
Audio-Visual Encoder (AVE) Talking head generation relies on strong correlations between phonemes and orofacial kinematics. While existing audio encoders (e.g., automatic speech recognition (ASR)-focused models) prioritize semantic extraction of speech signals, they fail to capture fine-grained mouth movement features critical for articulatory dynamics. To address this, we adopt SyncTalk’s pre-trained audio-visual encoder (Peng et al. 2024), adversarially trained on the LRS2 dataset (Afouras et al. 2018) with a lip-sync discriminator. This model extracts phoneme-aware mouth movement features $a$, establishing cross-modal alignment between speech signals and articulatory motion. These features are injected into the deformation field via attention-based modulation, controlling lip-related Gaussian primitive deformations.
Facial Action Encoder (FAE) Facial expressions during speech production involve complex muscle activations that extend beyond phoneme articulation, including subtle movements such as squinting, eyebrow elevation, and frowning. Current approaches exhibit limitations in modeling these dynamics. For example, some efforts regulate basic blinking patterns (Li et al. 2023; Tang et al. 2022; Liu and Hao 2023) but fail to capture nuanced facial muscle coordination. Blendshape-driven techniques (Peng et al. 2024) enable broader expression control through 3D shape approximations but lack direct correspondence to biomechanical muscle activations. To bridge this gap, we propose a facial action encoder grounded in the Facial Action Coding System (FACS) (Ekman and Friesen 1978). Our framework parameterizes facial dynamics using six anatomically defined Action Units (AUs: 1, 4, 5, 6, 7, 45), selected for two key advantages.
@@
Audio-Visual Encoder. To assess the effectiveness of AVE, we conduct ablation experiments on datasets with constrained head rotations (``Lieu", ``Shaheen" and ``Jae-in"), where subtle articulatory dynamics are critical for accurate lip synchronization. Specifically, for AVE-equipped methods (our M2DAO-Talker approach and SyncTalk (Peng et al. 2024)): we replace AVE with DeepSpeech (Amodei et al. 2016) (DS) audio encoding. For DS-based methods (TalkingGaussian (Li et al. 2025)), we substitute DS with AVE encoding. We maintain the identical architectures otherwise for fair comparison, and evaluate their performance across PSNR, Sync-C, and LMD. Experiments in Table 5 demonstrate that AVE implementations outperform DS variants across all metrics, which displays AVE's advantage in capturing mouth movement dynamics over DS's semantic speech features. It is speculated that AVE's focus on articulatory features enhances visual-articulatory alignment. Significantly, our M2DAO-Talker method maintains performance parity with SyncTalk and TalkingGaussian even when using suboptimal DS encoding, demonstrating obvious architectural resilience and superiority.
Backbone AVE DS PSNR $\uparrow$ Sync-C $\uparrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 35.932 7.793 2.423
35.672 6.560 2.484
TalkingGaussian (Li et al. 2025) 32.687 6.442 2.484
32.841 7.485 2.363
SyncTalk (Peng et al. 2024) 35.075 7.522 2.503
34.878 6.162 2.510
@@
Backbone AU BS PSNR $\uparrow$ Sync-C $\uparrow$ LMD $\downarrow$
M2DAO-Talker (Ours) 34.475 7.757 2.636
34.333 7.714 2.672
TalkingGaussian (Li et al. 2025) 32.042 6.387 2.712
31.938 6.397 2.712
SyncTalk (Peng et al. 2024) 33.902 7.671 2.685
33.905 7.675 2.660
@@
Facial Action Encoder. To evaluate the efficacy of FAE, we conduct cross-paradigm ablation experiments by substituting Action Units (AUs) and Blendshapes (BS) across seven datasets with varying expression complexity. For AU-based methods (our M2DAO-Talker approach and TalkingGaussian (Li et al. 2025)), we replace AUs with BS. Conversely, for BS-based methods, we substitute BS with AUs for facial expression control. The performance comparison across PSNR, Sync-C, and LMD are shown in Table 6, revealing that AU-to-BS substitution causes consistent metric degradation on average (PSNR: --0.142 dB, LMD: $+$0.036). Overall, AUs outperforms BS in almost scenarios, particularly in complex and diverse expression scenarios (e.g., ``May":$+$0.61 dB, ``Obama1" :$+$0.29 dB and ``Obama2" :$+$0.19 dB). This validates AUs' superiority in modeling fine-grained muscle activations critical for subtle expressions.
@@
To facilitate a more comprehensive comparison with SyncTalk and TalkingGaussian, we select five speech videos exceeding four minutes in duration from the High-Definition Talking Face (HDTF) dataset (Zhang et al. 2021). Unlike the datasets used in the main paper, these videos feature substantial torso motion and large-scale head rotations during speech, thereby presenting challenging real-world conditions that stress model generalization and robustness. We adopt the same self-reconstruction setup as in earlier experiments, and visualize the results in Figure 7. The following analysis highlights qualitative differences across the three methods.